# Deep Pre-trained Time Series Features for Tree Species Classification in the Dutch Forest Inventory

### v1

@@ Abstract
National Forest Inventory (NFI)s serve as the primary source of forest information, providing crucial tree species distribution data for carbon storage estimation and biodiversity assessments. However, maintaining these inventories requires labor-intensive on-site campaigns by forestry experts to identify and document tree species. Remote sensing approaches, particularly when combined with machine learning, offer opportunities to update NFIs more frequently and at larger scales. While the use of Satellite Image Time Series (SITS) has proven effective for distinguishing tree species through seasonal canopy reflectance patterns, current approaches rely primarily on Random Forest (RF) classifiers with hand-designed features and phenology-based metrics. Using deep features from an available pre-trained remote sensing foundation models offers a complementary strategy. These pre-trained models leverage unannotated global data and are meant to used for general-purpose applications and can then be efficiently fine-tuned with smaller labeled datasets for specific classification tasks. This work systematically investigates how deep features improve tree species classification accuracy in the Netherlands with few annotated data (1,462 pure species plots). We evaluate this question using three tree species datasets of varying difficulty: First the unbalanced Dutch NFI data with 13 species classes, second, unbalanced with 7 species classes, and, third, a balanced NFI dataset with 7 classes used by a RF algorithm used by recent related work. Data-wise, we extracted time-series data from Sentinel-1 (S1), Sentinel-2 (S2) and European Centre for Medium-Range Weather Forecasts Re Analysis v5 (ERA5) satellites data (January-December 2020) and Shuttle Radar Topography Mission (SRTM) data using Google Earth Engine. Our results demonstrate that fine-tuning a publicly available remote sensing time series foundation model outperforms the current state-of-the-art in NFI classification in the Netherlands by a large margin of up to 10% across all datasets. This demonstrates that classic hand-defined harmonic features are too simple for this task and highlights the potential of using deep AI features for data-limited application like NFI classification. By leveraging openly available satellite data and pre-trained models, this approach significantly improves classification accuracy compared to traditional methods and can effectively complement existing forest inventory processes.
@@
National Forest Inventory , deep learning , pre-trained remote sensing foundation model , time series classification
@@
[first]organization=Wageningen University,addressline=Droevendaalsesteeg 3 , city=Wageninegn, postcode=6708 PB, country=The Netherlands
@@ 1 Introduction
Forests play a significant role in mitigating climate change, adopting disaster prevention strategies, and providing ecosystem services, including sequestering carbon dioxide (CO2), providing wood materials, and serving as a source of biodiversity (Tomppo et al., 2010; FAO, 2020; Francini et al., 2024). Tree species diversity improves productivity of materials and resistance to natural disturbance (Jactel et al., 2017). Therefore, monitoring and tracking records of forests including spatial distribution of tree species are necessary for a sustainable forest management. National Forest Inventory (NFI)s are the primary source of information for various purposes such as sustainable forest management, industry investment planning, biodiversity monitoring, and Greenhouse Gas accounting (Tomppo et al., 2010; Bonannella, 2024). The Paris Agreement also requires the submission of a national carbon inventory, including carbon removals and reductions from forest lands (UNFCCC, 2015). Spatial tree species distribution information plays an important role in NFIs for various applications such as carbon storage estimation, forest management, and biodiversity assessments (Hermosilla et al., 2022; Blickensdörfer et al., 2024). Additionally, detailed tree species information is essential for national reports to the Forest Resource Assessment (FRA) of the Food and Agriculture Organization (FAO) and Forest Europe (Schelhaas et al., 2014). One of the key challenges in managing forest inventory is frequency and scale. Traditional inventory methods are based on sample-based field measurements conducted every 5 to 10 years (Tomppo et al., 2010). Current climate change and land-use changes due to economic growth cannot be captured in a timely manner using these traditional methods (Bonannella, 2024). Remote sensing helps improve and update NFIs (Hermosilla et al., 2022; Francini et al., 2024). Satellite sensors capture information about Earth every few days to several weeks, allowing us to monitor Earth’s condition more frequently and extensively, even in remote areas. However, the global coverage and complex data volumes in remote sensing make human visual interpretation impractical for large-scale analysis. Machine learning techniques have achieved significant improvements in Earth observations, including tree species classification, particularly when using multiple timestamps of images over large areas (Blickensdörfer et al., 2024; Francini et al., 2024; Hermosilla et al., 2022). The RF algorithm is one of the most popular machine learning algorithms for tree species classification due to its robustness, interpretability, and ability to handle high-dimensional data (Breiman, 2001). However, RF models require well-designed input features engineering, and this selection of appropriate features is crucial for model performance (Heaton, 2016). These choices depend on domain knowledge and target area characteristics such as climate and tree species variety, and often fail to include all necessary features (Ahlswede et al., 2023). In the context of NFIs, several state-of-the-art machine learning models utilizing RF have emerged (Hermosilla et al., 2022; Blickensdörfer et al., 2024; Francini et al., 2024)). These current methods for tree species classification rely on country-specific knowledge for input features, parameter settings, and high-quality data. However, scaling these methods to other countries at a national level presents challenges due to cost constraints and data availability limitations. Furthermore, deep learning models, which generally require large datasets to enhance performance, have remained largely unexplored for national-level tree species classification, primarily due to the limited size of available training datasets. These constraints collectively hinder the transferability of existing methods across different regions or countries. Deep learning models such as transformer architectures (Vaswani et al., 2023) have been recently introduced for forest monitoring, including tree species classification. This adoption is driven by increasing interest in multimodal and time-series data fusion in Remote Sensing (RS), enabled by the availability of big data and advancements in deep learning models (Li et al., 2022). Deep learning models can capture complex patterns in input data and create deep features that can be used for downstream tasks with classifiers and regressors including RF (Basu et al., 2015). While recent studies have achieved success in regional-scale tree species classification using high-quality labeled data, significant gaps remain in large-scale classifications (Fassnacht et al., 2016) for NFIs due to limited labeled data availability and high computational cost for training. Freely available pre-trained models, trained on large unlabeled datasets containing millions of pixels or images, have emerged as powerful tools for various downstream tasks. These models can achieve comparable or superior accuracy to traditional state-of-the-art machine learning approaches through fine-tuning without computationally expensive pre-training (Bommasani et al., 2022). Self-supervised learning, where models are trained without labels, has gained particular attention in RS applications (Wang et al., 2022). Using self-supervised learning as a model backbone with fine-tuning on limited labeled data has demonstrated significant accuracy improvements (Yu et al., 2022), particularly in time-series analysis tasks, while requiring less inductive bias (Dosovitskiy et al., 2021). The rapid growth of both labeled and unlabeled datasets for RS (Gorelick et al., 2017; Ahlswede et al., 2023) has enabled the development of various pre-trained models for tasks including tree classification (Lu et al., 2024). However, research comparing performance between fine-tuned pre-trained models and traditional approaches for tree species classification in NFIs remains limited.
This paper evaluates the effectiveness of deep features from freely available fine-tuned pre-trained models compared to a recently published random forest approach by Francini et al. (2024), who set a new state-of-the-art for Tree Species Identification in the Netherlands through a RF fitted on median reflectances and harmonic features.
The paper is structured as follows: section˜2 details the three datasets from the Dutch Forest Inventory used in this study. Section˜3 details the deep features we fine-tuned together with the harmonic features, as proposed by prior work.
@@ 2 Data
We use three datasets from the Dutch National Forest Inventory (NFI) at different complexity levels that we will use to compare the proposed fine-tuned deep time series features with the RF by Francini et al. (2024) fitted on harmonic time series features, as detailled later in section˜3.
Complex & imbalanced (COMB)
Simplified & imbalanced (SIMB)
Simplified & balanced (SIBA)
7 species classes, a balanced class distribution, by Francini et al. (2024). In total 13,790 samples and contains additional samples augmented by photointerpretation.
Table˜1 shows a detailed list of the tree species across all three datasets.
@@ 2.2 National Forest Inventory
Label Data Origin Ground truth data were collected manually through field measurements at 3,062 plots for the Dutch National Forest Inventory 6 between 2012 and 2013. Each plot contained a circular area with a variable radius (5 to 20 m) to ensure inclusion of at least 20 trees (Schelhaas et al., 2022). These plots correspond to 1 to 16 pixels at a $10\times 10$m resolution of Sentinel-1 and Sentinel-2. Due to privacy considerations, the precise coordinates of the plots were obtained under a confidentiality agreement with the Dutch government, with the requirement that the data be discarded after project completion. Prior to data disposal, we utilized the plot center coordinates from the total 3,062 pixel-level data points to extract satellite data at $10\times 10$ m pixel resolution from Google Earth Engine (GEE) (Gorelick et al., 2017).
@@
Class Aggregation The original NFI data contains 19 dominant tree species classes. However, we noticed that six dominant species classes have less than 10 samples and we decided to remove these classes to not extenstively rely in our evaluation on underrepresented tree species, similar to Kang et al. (2017). This resulted in the dataset Complex & Imbalanced (COMB) with 1,462 samples with 13 classes for the dominant species classification task. For the Simplified & Imbalanced (SIMB), we further aggregated the classes into the 7 tree species identical to Francini et al. (2024). This resulted in 1479 data samples.
@@
Finally, we also acquired the identical dataset used by Francini et al. (2024) which contains 13,790 data points evenly distributed across the aggregated classes (1,970 points per class). This dataset originated from the same NFI data but was augmented with additional labeled data points through visual interpretation of satellite imagery. We call Simplified & Balanced (SIBA) compared to aforementioned ones, as it contains 7 tree species classes and an even number of samples per class.
@@
2.3 Remote Sensing Time Series

### v2

# Assessing the Effectiveness of Deep Embeddings for Tree Species Classification in the Dutch Forest Inventory

@@ Abstract
National Forest Inventory (NFI) serves as the primary source of forest information, however, maintaining these inventories requires labor-intensive on-site campaigns by forestry experts to identify and document tree species. Embeddings from deep pre-trained remote sensing models offer new opportunities to update NFIs more frequently and at larger scales. While training new deep learning models on few data points remains challenging, we show that using pre-computed embeddings can directly outperform existing hand-designed features from Satellite Image Time Series (SITS) that have proven effective for distinguishing tree species through seasonal canopy reflectance patternsin combination with Random Forest (RF) classifiers. This work systematically investigates how deep embeddings improve tree species classification accuracy in the Netherlands with few annotated data (1,462 pure species plots). We evaluate this question on three embedding models: Presto, Alpha Earth, and Tessera, using three tree species datasets of varying difficulty. Data-wise, we compare the available embeddings from Alpha Earth and Tessera with dynamically calculated embeddings from a pre-trained Presto model, for which we extracted time series from Sentinel-1 (S-1), Sentinel-2 (S-2), and weather data, along with elevation data downloaded from Google Earth Engine. Our results demonstrate that fine-tuning a publicly available remote sensing time series pre-trained model outperforms the current state-of-the-art in NFI classification in the Netherlands, yielding performance gains of approximately 2–9 percentage points across datasets and evaluation metrics. This indicates that classic hand-defined features are too simple for this task and highlights the potential of using deep embeddings for data-limited applications such as NFI classification. By leveraging openly available satellite data and deep embeddings from pre-trained models, this approach significantly improves classification accuracy compared to traditional methods and can effectively complement existing forest inventory processes.
@@
National Forest Inventory , deep embeddings , pre-trained remote sensing model , time series classification
@@
[first]organization=Wageningen University,addressline=Droevendaalsesteeg 3 , city=Wageningen, postcode=6708 PB, country=The Netherlands
@@
[…]
[fourth]organization=University of Bonn,addressline=Meckenheimer Allee 174, city=Bonn, postcode=53115, country=Germany
@@ 1 Introduction
Forests provide critical climate-mitigation and ecosystem services, and tree-species diversity increases productivity and resistance to disturbances (Tomppo et al., 2010; FAO, 2020; Francini et al., 2024; Jactel et al., 2017). As a result, timely spatially explicit information on tree-species distribution is increasingly needed to support sustainable forest management, biodiversity assessments and carbon accounting (Tomppo et al., 2010; Bonannella, 2024; UNFCCC, 2015; Schelhaas et al., 2014; Hermosilla et al., 2022; Blickensdörfer et al., 2024). National Forest Inventory (NFI)s are the primary national reference for these applications, but they are fundamentally constrained by sampling-based field measurements typically repeated every 5–10 years, limiting their ability to capture rapid climate- and land-use-driven change at relevant scales (Tomppo et al., 2010; Bonannella, 2024). Spatial tree species distribution information plays an important role in NFIs for various applications such as carbon storage estimation, forest management, and biodiversity assessments (Hermosilla et al., 2022; Blickensdörfer et al., 2024). Additionally, detailed tree species information is essential for national reports to the Forest Resource Assessment (FRA) of the Food and Agriculture Organization (FAO) and Forest Europe (Schelhaas et al., 2014). Remote sensing has therefore become a key component of enhanced NFIs providing frequent, wall-to-wall observations that can complement plot networks and enable more up-to-date mapping (Hermosilla et al., 2022; Francini et al., 2024; White et al., 2025). Yet the scale and complexity of satellite time series make manual interpretation infeasible, motivating automated approaches. Machine learning methods, especially those leveraging multi-temporal imagery, have improved performance in Earth observation applications including tree-species classification (Blickensdörfer et al., 2024; Francini et al., 2024; Hermosilla et al., 2022). In operational NFI workflows, this has largely been addressed with Random Forest (RF) models over engineered time-series predictors, which perform well but hinge on careful feature design and local domain knowledge. The Random Forest algorithm is one of the most popular machine learning models for tree species classification due to its robustness, interpretability, and ability to handle high-dimensional data (Breiman, 2001). However, RF models require well-designed feature engineering, and the selection of appropriate features is crucial for model performance (Heaton, 2016). These choices depend on domain knowledge and target area characteristics such as climate and tree species variety, and often fail to include all necessary features (Ahlswede et al., 2023).
@@
In the context of NFIs, several state-of-the-art machine learning models utilizing RF have emerged (Hermosilla et al., 2022; Blickensdörfer et al., 2024; Francini et al., 2024)).The recent models have taken advantage of Satellite Image Time Series (SITS) to capture phenology differences among tree species (Hemmerling et al., 2021; Grabska and Sulikowska, 2025). These current methods for tree species classification rely on country-specific knowledge for selection and calculation of input features. Even though those models have achieved promising results, generalization and improvement of input data for model training remains the question. Deep learning models such as transformer architectures (Vaswani et al., 2023) have been recently introduced for forest monitoring, including tree species classification (Ma et al., 2024). This adoption is driven by increasing interest in multimodal and time-series data fusion in Remote Sensing (RS), enabled by the availability of big data and advancements in deep learning models (Li et al., 2022). Earth embeddings (Klemmer et al., 2025) from pre-computed features of deep learning models can capture complex patterns in in an easy-to-access form that can be directly used for downstream tasks with classifiers and regressors including RF (Basu et al., 2015). It is evident that models such as the temporal transformer have been introduced to capture complex temporal dependencies (Rußwurm_multi-temporal-land-classification_2018; Sainte Fare Garnot et al., 2020). While recent studies have achieved success in regional-scale tree species classification using high-quality labeled data, significant gaps remain in large-scale classifications (Fassnacht et al., 2016) for NFIs due to limited labeled data availability and high computational cost for training. Freely available pre-trained models and derived embeddings, trained on large unlabeled datasets containing millions of pixels or images, have emerged as powerful tools for various downstream tasks (Tseng et al., 2024; Feng et al., 2025; Brown et al., 2025). These models and embeddings can achieve comparable or superior accuracy to traditional state-of-the-art machine learning approaches through fine-tuning without computationally expensive pre-training (Bommasani et al., 2022). Self-supervised learning, where models are trained without labels, has gained particular attention in RS applications (Wang et al., 2022). Using self-supervised learning as a model backbone with fine-tuning on limited labeled data has demonstrated significant accuracy improvements (Yu et al., 2022), particularly in time-series analysis tasks, while requiring less inductive bias (Dosovitskiy et al., 2021). The rapid growth of both labeled and unlabeled datasets for RS (Gorelick et al., 2017; Ahlswede et al., 2023) has enabled the development of various pre-trained models for tasks including tree classification (Lu et al., 2024).
Despite rapid progress in machine learning for remote sensing, a key scientific challenge remains: under limited labeled data, can deep feature embedding representations capture spatial and temporal detail required by national forest inventories better than existing hand-crafted models. While existing approaches, such as country-specific RF models, depend heavily on expert-designed features and do not easily transfer to other ecological or sensing conditionsnew pre-trained embedding models promise scalable feature extraction but have not yet been systematically assessed for operational forest inventory settings.
In short, this study asks whether representations learned from large, unlabeled remote sensing datasets can outperform or complement traditional hand-crafted features in NFI workflows. Our main contributions are threefold:
  1. 1.
    We identify the lack of transferable, data-efficient tree species mapping methods as a limiting factor for timely and scalable forest monitoring.
  2. 2.
    We integrate deep embeddings from fine-tuned self-supervised models into a tree species classification pipeline, directly comparable to a state-of-the-art RF approach (Francini et al., 2024).
  3. 3.
    Using Dutch National Forest Inventory data, we conduct a systematic comparison between harmonic + medoid features and fine-tuned deep embeddings representations, analyzing accuracy, robustness, and generalization potential.
2 Methods
In this study, we fundamentally study the impact of time series feature embeddings on national forest inventory (NFI) by comparing hand-designed features from Francini et al. (2024), which represent the current state-of-the-art for Dutch NFI mapping, with deep learning embeddings from Presto, Tessera, and AlphaEarth. Table˜1 provides a high-level comparison in terms of data and usability criteria and the following-subsections summarize the respective models. Time series input is relevant to capture phenological sequential patterns, downloadable embeddings make applying deep embeddingsvery user-friendly and applicable for large-scale mapping tasks, and an open-source available fine-tunable model allow further modifications and possibly better downstream accuracy. Hence, we outline the methodological details of both complementary approaches in the next two sections before detailing the experimental setup in section˜4.
Table 1: Comparison of deep learning embedding models Presto (Tseng et al., 2024), TESSERA (Feng et al., 2025), and AlphaEarth (Brown et al., 2025) with traditional features (Francini et al., 2024) across time data and usability criteria relevant to NFI applications. Additional information includes encoder parameter count, pretraining dataset size, and pretraining type.
Model Time Series Input Open
Embeddings
Open
Model
Encoder Parameters Pretraining Data Size Pretraining Type
Presto (Tseng et al., 2024) 12 fixed monthly steps No Yes 0.4M 21.5M pixels Pixel-level
TESSERA (Feng et al., 2025) 40 random steps/year Yes (on request) Yes (on request) 40M 0.8B pixels Pixel-level
AlphaEarth (Brown et al., 2025) No (implicit temporal summarization) Yes No 480M–1B 3B images Image-level
Harmonic + Medoid features (Francini et al., 2024) Harmonic coefficients and seasonal median No N/A N/A N/A N/A
2.1 Deep Embeddings
These models reflect a recent development toward making pre-trained deep embeddings openly available (open embeddings) either directly for download, or through open-source, pre-trained, and fine-tunable models (open model).
2.1.1 Presto “Pretrained Remote Sensing Transformer” (Tseng et al., 2024)
Architecture and Training: Presto is a lightweight pixel-based Transformer trained on explicit multi-sensor time series. Each pixel is represented by a 12-month sequence of Sentinel-1 (S-1), Sentinel-2 (S-2), NDVI, ERA5 climate, and Dynamic World land-cover data, augmented with static contextual variables such as SRTM elevation and geographic coordinates. Presto learns through self-supervised masked autoencoding, reconstructing missing timesteps and sensor channels to capture temporal dependencies and handle missing data.
Suitability for NFI: Because Presto directly models monthly vegetation trajectories and integrates environmental context, it is highly suited for national forest inventory (NFI) tasks where phenological signals (e.g., seasonal greenness, moisture, canopy change) dominate. It can characterize both intra-annual and inter-annual forest dynamics. The trade-off is that Presto must be run locally, requiring access to raw satellite data, but it offers the strongest temporal interpretability for NFI-type monitoring.
Accessibility of model and embeddings: The Presto model is openly available (e.g., via GitHub) for download and local execution. However, pre-computed embeddings are not published for direct download — users must compute embeddings themselves from raw time-series inputs. This implies a heavier setup but full control over fine-tuning and feature extraction. We detail our setup in the section˜A.1.
2.1.2 Tessera “Temporal Embeddings of Surface Spectra” (Feng et al., 2025)
Architecture and Training: Tessera uses a dual-encoder Transformer architecture—one branch for S-1 SAR and another for S-2 optical data—processing entire annual time series per pixel (timestep $\times$ channels). Its self-supervised pretraining applies a modified Barlow Twins loss that enforces invariance between partial temporal views of the same pixel, encouraging the model to learn phenologically stable annual representations. The output is a 128-dimensional annual embedding per 10 m pixel, globally covering 2017–2024, available as pre-computed data on request.
Suitability for NFI: Tessera is well suited for NFI applications because it directly encodes annual spectral and backscatter evolution—the key seasonal information driving forest structure and productivity estimates—while being readily accessible as pre-computed embeddings. However, since it outputs annual summaries rather than continuous time series, it provides a coarser representation of intra-annual phenology compared to Presto.
Accessibility of model and embeddings: The Tessera repository is publicly accessible (MIT license) and supports inference and embedding extraction. Pre-computed embeddings are provided via the accompanying GeoTessera library for many years and regions, subject to request. Fine-tuning the model is possible since the code and architecture are open, but the underlying model weights need to be requested.
2.1.3 AlphaEarth Foundations “Embedding Field Model” (Brown et al., 2025)
Architecture and Training: AlphaEarth integrates diverse Earth observation sources (S-1/2, Landsat, GEDI, ERA5-Land, GRACE, and others) through a Space-Time Precision (STP) encoder that combines spatial and temporal attention with convolutional operations. It is trained as a multi-modal autoencoder with implicit decoders that reconstruct missing observations across time and modalities, regularized by a teacher–student consistency and contrastive alignment loss. The model learns continuous-time latent fields, producing annual “embedding layers” distributed globally via Google Earth Engine (Gorelick et al., 2017).
Suitability for NFI: While AlphaEarth embeddings reflect long-term temporal consistency and cross-sensor coherence, they do not directly input dense time series per pixel. Instead, they summarize information over broader “valid periods,” emphasizing spatial and modal fusion rather than explicit phenological sequencing. This makes AlphaEarth moderately suited for NFI applications—excellent for stable, multi-sensor mapping of forest attributes, but less precise for capturing fine-scale seasonal forest dynamics critical to phenology-based inventories.
Accessibility of model and embeddings: The embedding dataset of AlphaEarth (annual 10 m resolution embeddings) is openly accessible through the Earth Engine Data Catalog. However, the full trained model weights (for fine-tuning or re-training) are not officially released in full form by the authors. Therefore, end users can readily apply the embeddings, but cannot fine-tune the original model out of the box.
2.2 Harmonic and Seasonal Features
We reproduced closely the feature engineering approach of Francini et al. (2024), who published a new state-of-the-art classification approach for tree species mapping in the Netherlands. Their approach generate 209 features containing Seasonal Medoid Statistics and Harmonic Features that we describe in the next two paragraphs. While Francini et al. (2024) used only features extracted from S-2 time series, we found experimentally (detailed later in section˜5.1 and table˜3) that including features from S-1 further slightly improves the results. Since we will also use S-1 to extract deep embeddings (section˜2.1), we include S-1 features to ensure a fair comparison.
76 Seasonal Medoid Statistics
In total, 12 bands from S-1 and S-2 are selected and additional 7 indices are calculated from S-2 including Normalized Difference Vegetation Index (NDVI), Normalized Burn Ratio (NBR), Enhanced Vegetation Index (EVI), and Tasseled Cap transformations: Brightness (TCB), Wetness (TCW), Greenness (TCG), Angle (TCA). This results in 19 bands, where for each band the median band value in each of the 4 seasons, winter (January, February, December), spring (March-May), summer (June-August), and autumn (September-November), the medoids values are extracted. These features describe the data through season-wise statistics, but do not capture the dynamic change in seasonality itself.
133 Harmonic Features
To capture seasonal changes, sine and cosine harmonics
This approach was proposed by Francini et al. (2024) based on S-2 features. We added of S-1 features, as we found them to be slightly beneficial in accuracy (see section˜5.1).
@@ 3.2 National Forest Inventory
We use three datasets from the Dutch National Forest Inventory (NFI) at different complexity levels that we will use to compare the proposed deep embeddings with the RF by Francini et al. (2024) fitted on harmonic time series features, as detailed later in section˜2. Despite the relatively small overall forest area in the Netherlands (approximately 10–11 % of land cover), forests display spatial variability in species composition, age structure, and mixed stand types, reflecting responses to site conditions, management history, and ecological processes (Schelhaas et al., 2014).
Complex & imbalanced (COMB)
Simplified & imbalanced (SIMB)
Simplified & balanced (SIBA)
7 species classes, a balanced class distribution, as proposed by Francini et al. (2024). In total 13,790 samples and contains additional samples augmented by photointerpretation from Dutch NFI samples.
Orginal Dutch NFI samples (COMB and SIMB) are derived by random sampling with a density of 1 point per 100 ha(Schelhaas et al., 2014), which means that each random point is designated per square kilometer, and SIBA removed adjacent pixels with minimum distance among sample pixels was set to 15m, which reduced issues associated with spatial autocorrelation(Francini et al., 2024; Hermosilla et al., 2022). Table˜2 shows a detailed list of the tree species across all three datasets.
Label Data Origin Ground truth data were collected manually through field measurements at 3,062 plots for the Dutch National Forest Inventory 6 between 2012 and 2013. Each plot contained a circular area with a variable radius (5 to 20 m) to ensure inclusion of at least 20 trees (Schelhaas et al., 2022). These plots correspond to 1 to 16 pixels at a $10\times 10$m resolution of S-1 and S-2. Due to privacy considerations, the precise coordinates of the plots were obtained under a confidentiality agreement with the Dutch government, with the requirement that the data be discarded after project completion. Prior to data disposal, we utilized the plot center coordinates from the total 3,062 pixel-level data points to extract satellite data at $10\times 10$ m pixel resolution from Google Earth Engine (GEE) (Gorelick et al., 2017).
@@
Class Aggregation The original NFI data contains 19 dominant tree species classes. However, we noticed that six dominant species classes have fewer than 10 samples and we decided to remove these classes to not extensively rely on underrepresented tree species in our evaluation, similar to Kang et al. (2017). This resulted in the dataset Complex & Imbalanced (COMB) with 1,462 samples with 13 classes for the dominant species classification task. For the Simplified & Imbalanced (SIMB), we further aggregated the classes into the 7 tree species identical to Francini et al. (2024). This resulted in 1479 data samples.
@@
Finally, we also acquired the identical dataset used by Francini et al. (2024) which contains 13,790 data points evenly distributed across the aggregated classes (1,970 points per class). This dataset originated from the same NFI data but was augmented with additional labeled data points through visual interpretation of satellite imagery. We call Simplified & Balanced (SIBA) compared to aforementioned ones, as it contains 7 tree species classes and an even number of samples per class. To augment the original Dutch NFI plots, the authors performed a rigorous three-step validation process: (1) visual photointerpretation using high-resolution orthophotos and Google satellite imagery to remove plots affected by land-cover change or mixed-species compositions; (2) stand-polygon identification to extract homogenous spectral samples; and (3) spatial thinning, where a minimum distance of 15 m was maintained between pixels to minimize spatial autocorrelation and label noise. The final dataset comprises 13,790 samples (1,970 per class).
@@
3.3 Remote Sensing Time Series for Presto and harmonic + medoid calculation